# Rethinking the Atmospheric Scattering-driven Attention via Channel and Gamma Correction Priors for Low-Light Image Enhancement

### v3

@@ Abstract
Enhancing low-light images remains a critical challenge in computer vision, as does designing lightweight models for edge devices that can handle the computational demands of deep learning. In this article, we introduce an extended version of the Channel-Prior and Gamma-Estimation Network (CPGA-Net), termed CPGA-Net+, which incorporates an attention mechanism driven by a reformulated Atmospheric Scattering Model and effectively addresses both global and local image processing through Plug-in Attention with gamma correction. These innovations enable CPGA-Net+ to achieve superior performance on image enhancement tasks for supervised and unsupervised learning, surpassing lightweight state-of-the-art methods with high efficiency. Furthermore, we provide a theoretical analysis showing that our approach inherently decomposes the enhancement process into restoration and lightening stages, aligning with the fundamental image degradation model. To further optimize efficiency, we introduce a block simplification technique that reduces computational costs by more than two-thirds. Experimental results validate the effectiveness of CPGA-Net+ and highlight its potential for applications in resource-constrained environments.
@@
Atmospheric Scattering Model, Low-Light Image Enhancement, Gamma Correction, Channel Prior, Explainable AI.
@@ I Introduction
Low-light image capture, whether indoors or outdoors, poses significant challenges for accurate visual analysis. The limited light reflection often results in degraded image quality, including color inaccuracies and increased noise levels. These issues can significantly affect the performance and reliability of light-sensitive applications, such as transportation surveillance and Advanced Driver Assistance Systems. Therefore, it is crucial to address these challenges to ensure the effective operation of systems under low-light conditions.
The problems of low-light image enhancement (LLIE) are commonly addressed using two main methods: Histogram Equalization [1] and Retinex [2]. Histogram Equalization works by enhancing contrast through the redistribution of grayscale values. On the other hand, the Retinex theory divides the image into reflectance and illumination components to improve reflectance and overall image quality. Techniques such as Single Scale Retinex [3] and Multi-Scale Retinex [4] are particularly effective in preserving details and managing complex lighting conditions.
With technological advancements, various deep learning-based methods [5, 6, 7] have been proposed to enhance the quality of low-light images. However, these methods often require substantial computational resources, which limits their practical application on real-world devices. Therefore, designing lightweight and efficient image enhancement techniques is crucial. In our previous study, we introduced the CPGA-Net [8], which combines Retinex theory with the Atmospheric Scattering Model (ATSM) [9] and utilizes gamma correction for both global and local processing; it highlights the importance of gamma correction in LLIE. CPGA-DIA [10] explores exposure correction and low-light enhancement issues through dynamic gamma adjustment, showing that gamma correction can be efficient and effective for the enhancement process even in deep learning frameworks.
@@
Building on these ideas with a theoretical-based structure and gamma-correction prior, we expanded CPGA-Net by rethinking the theoretical equation as an attention mechanism and transitioning the gamma estimation module from global processing to local processing. Our main contributions are as follows:
  • Extended CPGA-Net: We propose an enhanced version of the Channel-Prior and Gamma-Estimation Neural Network (CPGA-Net) called CPGA-Net+. This model achieves state-of-the-art (SOTA) image quality and efficiency performance for both supervised and unsupervised learning, making it a lightweight and practical solution for real-world applications, as shown in Fig. 1.
  • Theoretical-based Attention for Illumination: We modularized the Atmospheric Scattering Model into a block design and incorporated gamma correction into the local branch. This architecture modification significantly improves image structure and detail, maximizing the efficiency of prior knowledge for illumination to strengthen the overall image quality.
  • See Through the Mechanism Behind the Image Enhancement: We evaluate our approach through the degradation formula and optimization, demonstrating that the optimal solutions align with the output features from the neural network, ensuring strong interpretability and robustness.
  • Block Design Simplification for Lightweight Implementation: We can directly remove the local branch of our network during inference, which induces a slight performance drop but significantly improves efficiency, reducing the number of parameters and FLOPS by about 66.8% and 77.1%, respectively.
@@ II-A Deep Learning-Based LLIE
With the continuous development of LLIE, Retinex theory has increasingly demonstrated its potential in conjunction with deep learning techniques. Several methods based on this approach address low-light environments. For instance, while LIME [11] differs from directly decomposing images according to Retinex theory, it primarily relies on estimating the illumination map of low-light images for enhancement. RetinexNet [5] and KinD [7] decompose images into reflectance and illumination components during decomposition. In the adjustment phase, they adjust the illumination component’s brightness and denoise the reflectance component, ultimately merging them based on the theory to restore natural images. EnlightenGAN [6] proposes an unsupervised Generative Adversarial Network (GAN) that can be trained without paired low/normal light images. It introduces a global-local discriminator structure, self-regularized perceptual loss fusion, and attention mechanisms to enhance image quality. LLFlow [12] presents a flow-based approach for LLIE by modeling distributions of normally exposed images. It improves traditional methods by using an illumination-invariant color map as the prior distribution rather than a Gaussian distribution. The process features an encoder to extract stable color attributes and an invertible network to map low-light images to normally exposed image distributions, aiming for improved enhancement performance.
@@ II-B Lightweight LLIE
In the context of LLIE, developing lightweight methods is crucial for practical deployment, often requiring sophisticated techniques to achieve both efficiency and effectiveness. For example, Zero-DCE [13] replaces the direct image enhancement process with a curve-fitting approach, introducing a series of reference-free loss functions that reduce the computational burden, achieving an efficient and lightweight design. RUAS [14] builds upon Retinex theory by proposing a Retinex-inspired model that leverages prior information from low-light images, combined with a distillation unit-based search architecture and a cooperative bilevel search strategy, maintaining high performance while achieving a lightweight design. IAT [15] decomposes the task into local and global processing components. The local branch leverages a convolution-based Transformer to perform image restoration and enhancement. In contrast, the global branch utilizes global priors, including color transformation matrices and gamma correction, to apply global adjustments across different exposure conditions, thereby attaining efficient and lightweight performance improvements. PairLIE [16] deviates from the traditional Retinex approach of directly decomposing images; instead, it removes noise through a self-supervised mechanism before decomposition. It shows that training on low-light images of the same scene with different exposures better learns features. Finally, it merges them using a simple convolutional network to achieve a lightweight design. Inspired by Retinex theory and ISP (Image Signal Processor) frameworks, FLIGHT-Net [17] features Scene Dependent Illumination Adjustment for illumination and gain processing and Global ISP Network Block for compact color correction and denoising. This design optimizes for both efficiency and lightweight operation.
@@ II-C Insights and Innovations
Our proposed method builds on two key insights from prior work: theory-driven attention mechanisms and lightweight design principles in LLIE. Our approach leverages the theoretical basis of perceiving visual information through air turbulence—a principle effectively utilized in Retinex-based [5, 6, 7, 11, 12, 14, 16] and ATSM methods [8, 10, 18]. We incorporate this concept into a streamlined attention module that selectively enhances key features, guided by prior knowledge [8, 13, 19], to improve detail preservation and contrast. Furthermore, following [8, 10, 15], gamma correction is integrated into our model and has extended its application from global to local processing to maximize adaptability via environmental characteristics. These attention modules and formulae turn our model apart by integrating theoretical principles with practical design, delivering high-quality enhancement while preserving a lightweight architecture.
@@ III Methodology
Figure 2: The schematic of our proposed approach, CPGA-Net+.
III-A The Connection Between Theoretical Equations and Low-light Image Enhancement
Guo et al. [11] figured out that the Retinex theory [2] and the ATSM [9] have a strong relationship that can represent each other by reformulating their equations. Retinex theory assumes that the received image can be decomposed into illumination and reflectance components, as shown in Eq. (1):
$S=i\cdot R$ (1)
@@ III-A The Connection Between Theoretical Equations and Low-light Image Enhancement
Based on Dong et al. [18], the low light image $I$ can be seen as $1-L$, where $L$ represents the low-light image, and $J$ can be seen as $1-R$, where $R$ reflects the important characteristics of the input image. The above substitutions are performed under the normalized pixel values in [0, 1]. With these substitutions, we rewrite Eq. (2) into the following form:
@@
where $\tilde{A}=1-A$ and $\tilde{t}=1/t$. The model Eq. (3) described is the cornerstone of our neural network design. Part of the information of the reflectance $R$ comes from the known image $L$, part of it comes from an unknown image $\tilde{A}$, and the proportion sum of their contributions is limited to 1. When $L$ is very dark or noisy (the scene information is less reliable), the contribution of $L$ is lowered, and the contribution of $\tilde{A}$ is increased; when $L$ is relatively bright and less noisy (the scene information is more reliable), the contribution of $L$ is increased, and the contribution of $\tilde{A}$ is reduced. So, $\tilde{t}$ should reflect the intensity level of $L$ in some way.
@@
This reformulation reveals an alternative imaging perspective, where $L$ is linked to the characteristic of atmospheric light, which predominantly includes environmental interference, $\tilde{A}$ corresponds to the unknown noise-free image, and $R$ corresponds to the reflectance in a linear relationship between $L$ and $\tilde{A}$. The underlying mechanism of this formulation matches the phenomenon of our atmospheric scattering-based approach and will be discussed in Section III-D.
In our previous work, CPGA-Net [8], we successfully utilized these theoretical equations in deep learning form. In this article, we extended the idea and proposed an attention mechanism called the Channel-Prior block (CP block), which modularizes the relationship of both equations into a systematic form. We restructured the module with convolutions and a ResBlock while fusing features with the original RGB channels at each step to streamline the module into an attention-block design. This helps in activating the feature map to align with the original channels. In CPGA-Net, three-channel priors are selected as the input for $\tilde{t}$ estimation: the Bright Channel Prior (BCP), the Dark Channel Prior (DCP), and the luminance channel (Y component from the YCbCr color space). They can be defined respectively as:
$I^{\text{bright}}=\max_{c\in\{r,g,b\}}\left(I^{c}\right)$ (4)
$I^{\text{dark}}=\min_{c\in\{r,g,b\}}\left(I^{c}\right)$ (5)
$I^{\text{luminance}}=0.299\cdot I^{r}+0.584\cdot I^{g}+0.114\cdot I^{b}$ (6)
where $I^{c}$ represents the color channel $c$ of the input image $I$. These common features represent the brightness variation in different environments and are widely used in traditional methods [19]. The combination of channel priors shows sensitivity to contrast, which is an important clue in representing the atmospheric transmittance $\tilde{t}$ to guide the enhancement, as shown in Fig. 3.
Figure 3: The comparison between the brightness representation. From top to bottom are the original image, edge detection of the luminance channel $I^{\text{luminance}}$, and edge detection of the channel-prior which is $\max(I^{\text{bright}},I^{\text{dark}},I^{\text{luminance}})$.
Additionally, to transform the priors into a high-dimensional processing module rather than maintaining the original channels, we simplified the luminance channel into a more basic representation—specifically, the mean of the channels, akin to the concept of intensity (such as the I component in the HSI color space), which also indicates image brightness. Therefore, the channel priors are depicted as:
[…]
where $f$ is the input feature map, and $F^{\text{CP}}$ denotes the channel prior features. If $c$ consists of RGB channels, $f$ will equal the $I$ that appears in Eq. (4), (5), and (6) As a result, the channel priors are simplified to the input channel’s max, min, and mean, making the attention module more responsive to the overall contrast control of the image.
@@
For the $\tilde{A}$ estimation, which captures detailed features and reconstructs the image, we redesigned it as a mini-U-Net-based architecture with encoder and decoder pathways connected by skip connections. This design effectively captures and preserves spatial information at multiple scales, enhancing the model’s ability to reconstruct fine details in the image—a technique commonly used in LLIE, such as [5, 6, 7]. Due to considerations for lightweight efficiency, we only downsample the input once, reducing computational complexity while maintaining adequate feature extraction capabilities. This approach ensures that the model remains efficient and suitable for real-time applications or scenarios with limited computational resources without significantly compromising the quality of the reconstructed image.
@@
$R_{\text{att}}(f)=\tilde{t}(f,F^{\text{CP}})L^{\prime}(f)+\big{[}1-\tilde{t}(f,F^{\text{CP}})\big{]}\tilde{A}(f)$ (8)
where $L^{{}^{\prime}}$ is the mapped input tensor with a matching channel for formula calculation, $\tilde{t}(f,F^{\text{CP}})$ indicates the derived transmittance with input feature map $f$ and channel-prior features $F^{\text{CP}}$, and the structure is shown in Fig. 4.
Figure 4: The block diagram of the Channel-Prior Block (CP block), where c denotes a concatenation operation.
III-B Relinking the Gamma Correction to Global-Local Processing
Not only taking advantage of the characteristics of low-light images, CPGA-Net [8] also combines gamma correction based on IAT [15]. Gamma correction is a simple technique that adjusts all the pixels with pointwise exponential operations, as shown in Eq. (9):
$s=r^{\gamma}$ (9)
$\begin{split}\hat{R}&=\text{IAAF}(R^{\gamma},R)=(R\cup R^{\gamma})-(R\cap R^{\gamma})\\ &\approx R+R^{\gamma}-\cap(R,R^{\gamma})\end{split}$ (10)
where the enhanced image $\hat{R}$ is created by combining $R$ and $R^{\gamma}$ while removing any overlapping elements and $\cap(R,R^{\gamma})$ represents the intersection estimation for finding similar features across gamma-corrected and uncorrected images.
First, based on the findings from [20, 21], we know that downsampling the image resolution by a limited scale does not significantly impact performance while greatly reducing computational cost. Therefore, we redesigned the process of gamma value estimation, as shown in Fig. 2. We reduced the feature resolution in the first ResCBAM block by adjusting the stride from 1 to 2. This adjustment enables subsequent computations to process half-sized images while leveraging residual and downsampling connections. This design improves the learnability of the IAAF module, facilitating more effective integration of global and local processing.
Furthermore, we adapt this operation to function within the Channel-Prior Block by incorporating it directly into the attention process, allowing it to act as a global feature that estimates the optimal gamma-correction value for each feature channel rather than applying gamma correction only at the final step as commonly done in [8, 15]. The modified attention output can then be expressed as:
$\hat{R}_{\text{att}}=\text{IAAF}(R_{\text{att}}^{\gamma},R_{\text{att}})+R_{\text{att}}$ (11)

### v4

# Rethinking Theoretical Illumination for Efficient Low-Light Image Enhancement

@@ Abstract
Enhancing low-light images remains a critical challenge in computer vision, as does designing lightweight models for edge devices that can handle the computational demands of deep learning. This article introduces an extended version of the Channel-Prior and Gamma-Estimation Network (CPGA-Net), termed CPGA-Net+, incorporating the theoretically-based Attentions for illumination in local and global processing. Additionally, we assess our approach through a theoretical analysis of the block design by introducing both an ultra-lightweight and a stronger version, following the same design principles. The lightweight version significantly reduces computational costs by over two-thirds by utilizing the local branch as an auxiliary component. Meanwhile, the stronger version achieves an impressive balance by maximizing local and global processing capabilities. Our proposed methods have been validated as effective compared to recent lightweight approaches, offering superior performance and scalable solutions with limited computational resources. The source code will be released at https://github.com/Shyandram/CPGA-Net_Plus.
@@
Atmospheric Scattering Model, Low-Light Image Enhancement, Lightweight, Channel Prior, Explainable AI.
@@ I Introduction
Whether indoors or outdoors, low-light image capture poses significant challenges for accurate visual analysis. The limited light reflection often results in degraded image quality, including color inaccuracies and increased noise levels. These issues can significantly affect the performance and reliability of light-sensitive applications, such as transportation surveillance and Advanced Driver Assistance Systems. Therefore, it is crucial to address these challenges to ensure the effective operation of systems under low-light conditions.
The problems of low-light image enhancement (LLIE) are commonly addressed using two primary methods: Histogram Equalization [1] and Retinex [2]. Histogram Equalization works by enhancing contrast through the redistribution of grayscale values. On the other hand, the Retinex theory divides the image into reflectance and illumination components to improve reflectance and overall image quality. Techniques such as Single Scale Retinex [3] and Multi-Scale Retinex [4] are particularly effective in preserving details and managing complex lighting conditions.
With technological advancements, various deep learning-based methods [5, 6, 7] have been proposed to enhance the quality of low-light images. However, these methods often require substantial computational resources, which limits their practical application on real-world devices. Therefore, designing lightweight and efficient image enhancement techniques is crucial. In our previous study, we introduced the CPGA-Net [8], which combines Retinex theory with the Atmospheric Scattering Model (ATSM) [9] and utilizes gamma correction for both global and local processing; it highlights the importance of gamma correction in LLIE. CPGA-DIA [10] explores exposure correction and LLIE issues through dynamic gamma adjustment, showing that gamma correction can be efficient and effective for the enhancement process even in deep learning frameworks.
@@
In our in-depth analysis of our previous work, CPGA-Net [8], we observed a significant discrepancy between our initial hypothesis and its empirical behavior. Specifically, the local branch, intended to provide supportive features for the global processing, converged towards a near-identity mapping. This indicates that the local branch offers minimal contribution during inference, rendering it largely redundant in the final design. This critical finding, however, presents two compelling opportunities: leveraging this redundancy for ultra-lightweight structural pruning and unleashing the potential capacity by stimulating it for more powerful performance.
Acting on these insights, this paper introduces a significant evolution of the CPGA-Net framework by addressing these opportunities. We expanded the theoretical structure by transforming the theoretical equations into an attention mechanism, and we have conducted a thorough analysis of the underlying learning mechanisms and their potential. Our main contributions are as follows:
  • Extended CPGA-Net: We propose an enhanced version of the Channel-Prior and Gamma-Estimation Neural Network (CPGA-Net) called CPGA-Net+. This model achieves state-of-the-art (SOTA) image quality and efficiency performance for supervised and unsupervised learning. It is a lightweight and practical solutions for real-world applications, as shown in Fig. 1.
  • Theoretically-based Attention for Illumination: We modularized the Atmospheric Scattering Model into a block design and incorporated gamma correction into the local branch. This architecture modification significantly improves image structure and detail, maximizing the efficiency of prior knowledge for illumination to strengthen the overall image quality.
  • Insight-Driven Network Simplification and Refinement: We identify a paradoxical training dynamic within our baseline model, which allows us to propose both a novel, training-free pruning method for simplification and an advanced fusion module that intelligently re-balances the network for enhanced performance.
@@ II-A Deep Learning-Based LLIE
With the continuous development of LLIE, Retinex theory has increasingly demonstrated its potential in conjunction with deep learning techniques. Several methods based on this approach address low-light environments. For instance, while LIME [11] differs from directly decomposing images according to Retinex theory, it primarily relies on estimating the illumination map of low-light images for enhancement. RetinexNet [5] and KinD [7] decompose images into reflectance and illumination components during decomposition. In the adjustment phase, they adjust the illumination component’s brightness and denoise the reflectance component, ultimately merging them based on the theory to restore natural images. EnlightenGAN [6] proposes an unsupervised Generative Adversarial Network (GAN) that can be trained without paired low/normal light images. It introduces a global-local discriminator structure, self-regularized perceptual loss fusion, and attention mechanisms to enhance image quality. LLFlow [12] presents a flow-based framework for LLIE that uses an illumination-invariant color map instead of a Gaussian distribution. It features an encoder for stable color attribute extraction and an invertible neural network to transform low-light images into normally exposed distributions, improving image quality and visibility. LLFormer [13] integrates a specialized illumination attention module and adaptive exposure fusion, allowing it to dynamically adjust to varying lighting conditions. While these methods effectively addressed the LLIE problems, they overlooked crucial computational issues essential for practical implementation.
@@ II-B Lightweight LLIE
In the context of LLIE, developing lightweight methods is crucial for practical deployment, often requiring sophisticated techniques to achieve both efficiency and effectiveness. For example, Zero-DCE [14] replaces the direct image enhancement process with a curve-fitting approach, introducing a series of reference-free loss functions that reduce the computational burden, achieving an efficient and lightweight design. RUAS [15] builds upon Retinex theory by proposing a Retinex-inspired model that leverages prior information from low-light images, combined with a distillation unit-based search architecture and a cooperative bilevel search strategy, maintaining high performance while achieving a lightweight design. IAT [16] decomposes the task into local and global processing components. The local branch leverages a convolution-based Transformer to perform image restoration and enhancement. In contrast, the global branch utilizes global priors, including color transformation matrices and gamma correction, to apply global adjustments across different exposure conditions, thereby attaining efficient and lightweight performance improvements. PairLIE [17] deviates from the traditional Retinex approach of directly decomposing images; instead, it removes noise through a self-supervised mechanism before decomposition. It shows that training on low-light images of the same scene with different exposures learns features better. Finally, it merges them using a simple convolutional network to achieve a lightweight design. Inspired by Retinex theory and ISP (Image Signal Processor) frameworks, FLIGHT-Net [18] features Scene Dependent Illumination Adjustment for illumination and gain processing and Global ISP Network Block for compact color correction and denoising. This design optimizes for both efficiency and lightweight operation. LYT-Net [19] is a recent advancement in LLIE that enhances efficiency by operating in the YUV color space. It uses a lightweight convolutional module to process the luminance (Y) channel with a Transformer block while denoising the chrominance channels (U/V). This approach significantly reduces computational costs and illustrates how thoughtful design integrates Transformer modeling with real-time performance.
@@ II-C Insights and Innovations
Our method is inspired by two foundational insights from prior literature: the effectiveness of theory-guided attention mechanisms and the importance of lightweight design for practical LLIE. Drawing from principles observed in Retinex-based approaches [11, 5, 7] and atmospheric scattering models (ATSM)  [20, 8, 10], we reinterpret the interaction between light and image degradation—particularly the way visual information is distorted by ambient air turbulence or illumination imbalance—as a guiding prior for attention learning. This perspective informs the design of our attention module, which selectively emphasizes structurally and visually significant regions during enhancement. Unlike conventional attention mechanisms that rely solely on data-driven cues, our model incorporates domain-specific priors  [21, 8, 19] to guide feature modulation. This allows for better contrast enhancement and preservation of fine-grained details, even under challenging lighting conditions. Additionally, building upon the practices in [8, 16, 18], we integrate gamma correction not just as a global post-processing step, but as a learnable and spatially adaptive component within our network. This shift from global to localized gamma adjustment increases the model’s adaptability to varying exposure and illumination distributions across an image.
Overall, our architecture exemplifies the fusion of physical imaging principles with deep learning efficiency. By embedding theoretical insights into a compact and modular attention framework, our model achieves high-quality enhancement with minimal computational overhead—making it suitable for real-world low-light applications on edge devices.
@@ III Methodology
Figure 2: The schematic of our proposed approach, CPGA-Net+. This paper presents the baseline model CPGA-Net+ and two improvements: CPGA-Net+ (BDSF) for an ultra-lightweight design and CPGA-Net++ for enhanced performance, as detailed on the right.
III-A Preliminary
Our research is based on a key finding by Guo et al. [11], which indicates that a strong theoretical relationship exists between Retinex theory [2] and the Atmospheric Scattering Model (ATSM) [9]. This connection allows for a reinterpretation of LLIE through the lens of haze removal principles. Retinex theory assumes that the received image can be decomposed into illumination and reflectance components, as shown in Eq. (1):
$S=i\cdot R,$ (1)
@@ III-A Preliminary
Based on Dong et al. [20], the low-light image $I$ can be seen as 1L1-L, where $L$ represents the low-light image, and $J$ can be seen as 1R1-R, where $R$ reflects the important characteristics of the input image. The above substitutions are performed under the normalized pixel values in [0, 1]. With these substitutions, we rewrite Eq. (2) into the following form:
@@
where $\tilde{A}=1-A$ and $\tilde{t}=1/t$. The model described in Eq. (3) is the cornerstone of our neural network design. Part of the information of the reflectance $R$ comes from the known image $L$, and part of it comes from an unknown image $\tilde{A}$, and the proportion sum of their contributions is limited to 1. When $L$ is very dark or noisy (the scene information is less reliable), the contribution of $L$ is lowered, and the contribution of $\tilde{A}$ is increased; when $L$ is relatively bright and less noisy (the scene information is more reliable), the contribution of $L$ is increased, and the contribution of $\tilde{A}$ is reduced. So, $\tilde{t}$ should reflect the intensity level of $L$ in some way.
In our previous work, CPGA-Net [8], we successfully utilized these theoretical equations in deep learning form as the local processing in image enhancement. Moreover, to further refine LLIE, we also considered global processing technique, gamma correction, as the brightness control, which is a simple technique that adjusts all the pixels with pointwise exponential operations, as shown in Eq. (4):
$s=r^{\gamma},$ (4)
$\begin{split}\hat{R}&=\text{IAAF}(R^{\gamma},R)=(R\cup R^{\gamma})-(R\cap R^{\gamma})\\ &\approx R+R^{\gamma}-\cap(R,R^{\gamma}),\end{split}$ (5)
where the enhanced image $\hat{R}$ is created by combining $R$ and $R^{\gamma}$ while removing any overlapping elements and $\cap(R,R^{\gamma})$ represents the intersection estimation for finding similar features across gamma-corrected and uncorrected images. CPGA-Net enhances image quality and contrast with good perspective evaluation by combining local and global processing.
III-B The Connection Between Theoretical Equations and Low-Light Image Enhancement
In this section, we thoroughly evaluate the CPGA-Net from the block structure perspective, including reevaluating the application of prior knowledge as Atmospheric Scattering-driven Attention and the mixture of local-global processing as Plug-in Attention with Gamma Correction.
III-B1 Atmospheric Scattering-driven Attention
@@
First of all, the reformulation of Eq. (3) reveals an alternative imaging perspective, where L is linked to the characteristic of atmospheric light, which predominantly includes environmental interference, $\tilde{A}$ corresponds to the unknown noise-free image, and R corresponds to the reflectance in a linear relationship between $L$ and $\tilde{A}$. This formulation’s underlying mechanism matches the phenomenon of our atmospheric scattering-based approach, performing an insightful attention relationship of these factors. Further analysis will be discussed in the Appendix.
Built upon these relationships, we extended the local processing as an attention mechanism called the Channel-Prior block (CP block), which modularizes the relationship of underlying equations into a systematic form. We restructured the module with convolutions and a ResBlock [22], fusing features with the original RGB channels to streamline it into an attention-block design.
For the attention to the brightness correlated priors for transmittance $\tilde{t}$ estimation, we adopt channel-prior features $F^{CP}$, which are sensitive to contrast and brightness variations. These are derived from the input feature map $f$:
$F^{\text{CP}}(f)=\text{concat}\left(\max_{c}\left(f^{c}\right),\min_{c}\left(f^{c}\right),\underset{c}{\text{mean}}\left(f^{c}\right)\right),$ (6)
where $f$ is the input feature map, and $F^{\text{CP}}$ denotes the channel prior features, which include brightness sensitive channel priors: Bright Channel Prior (BCP), Dark Channel Prior (DCP), and the luminance channel (I component in the HSI color space). These priors have been widely used in prior research to capture the characteristics of low-light images. In our implementation, we simplify the prior into a high-dimensional representation by reducing the channel priors to the input channels’ maximum, minimum, and mean values. This approach improves the attention module’s responsiveness in controlling the image’s overall contrast.
@@
For the $\tilde{A}$ estimation, which captures detailed features and reconstructs the image, we redesigned it as a mini-U-Net-based architecture with encoder and decoder pathways connected by skip connections. This approach ensures that the model remains efficient and suitable for real-time applications or scenarios with limited computational resources without significantly compromising the quality of the reconstructed image.
@@
$R_{\text{att}}(f)=\tilde{t}(f,F^{\text{CP}})L^{\prime}(f)+\big{[}1-\tilde{t}(f,F^{\text{CP}})\big{]}\tilde{A}(f),$ (7)
where $L^{{}^{\prime}}$ is the mapped input tensor with a matching channel for formula calculation, $\tilde{t}(f,F^{\text{CP}})$ indicates the derived transmittance with input feature map $f$ and channel-prior features $F^{\text{CP}}$, and the entire structure is shown in Fig. 4.
Figure 4: The Channel-Prior Block (CP block) block diagram, where c denotes a concatenation operation.
III-B2 Plug-in Attention with Gamma Correction
As analyzed in [10], the adaptive gamma value serves as an environmental factor representing the overall illumination conditions, directly reflecting the restoration’s state. To leverage the prior information of environment representation, we have integrated a global processing operation within the Channel-Prior Block and incorporated it directly into the attention mechanism. This approach enables the global feature to estimate the optimal gamma-correction value for each feature channel, rather than limiting gamma correction to just the final step, as is often done in previous studies [8, 16]. The modified attention can then be expressed as:
$\hat{R}_{\text{att}}=\text{IAAF}(R_{\text{att}}^{\gamma},R_{\text{att}})+R_{\text{att}},$ (8)